A rug placed in a 10 ft x 12 ft room covers two-thirds of the floor area and leaves a uniform strip of bare floor around teh edges. Find the dimensions of the rug.
I dont need the answer, just the step-by-step process.
: A rug placed in a 10 ft x 12 ft room covers two-thirds of the floor area and leaves a uniform strip of bare floor around teh edges. Find the dimensions of the rug.
I dont need the answer, just the step-by-step process.

Answer by nerdybill(1129)

The important thing is to draw a diagram. Label all the information provided.
.
Let x = width of strip between the edge of the rug and wall
.
The room is 10 by 12 feet.
.
Therefore the rug must be:
(10-x) by (12-x)
.
The area of the rug is equals to 2/3 of the room:
(10-x)(12-x) = 2/3(10*12)
(10-x)(12-x) = 2/3(120)
(10-x)(12-x) = 2(40)
(10-x)(12-x) = 80
120-10x-12x+x^2 = 80
120-22x+x^2 = 80
x^2-22x+120 = 80
x^2-22x+40 = 0
(x-20)(x-2) = 0
.
x = {2,20}
.
Therefore the rug is:
10 feet by 8 feet